# Improved Belief Propagation Decoding Algorithms for Surface Codes

### v3

@@ IV.2.2 Toric code
For the serial schedule, all three decoding algorithms demonstrated error correction capabilities. Among those, BP-OTS performs better than MBP at low physical error rates, but its performance is surpassed by MBP when the physical error rate exceeds 7%. The performance of EWAInit-BP, on the other hand, consistently outperforms these two methods across all tested conditions, with the logical error rate at $L=6$ being comparable to that of the other two algorithms at $L=8$, and the intersection of the logical error rate curves is approximately 15%. In terms of iteration count, both EWAInit-BP and MBP converge in remarkably few iterations (often fewer than 5).
It is noteworthy that for the toric code, the performance of EWAInit-BP under parallel scheduling is not significantly different from serial scheduling. This is due to the method’s inherent ability to effectively break trapping sets, combined with the geometric symmetry of the Toric code, which allows for less dependency on the scheduling method. However, for the following topological codes with more complex stabilizer configurations, serial scheduling shows better performance.
IV.2.3 Planar surface code
The boundaries of the planar surface code are no longer contiguous, resulting in the emergence of some stabilizers with weight three and qubits that only connect two stabilizers at the boundaries. This configuration leads to a slight reduction in the error correction capabilities of the planar surface code compared to the Toric code. We compare the decoding logical error rates of LLRBP4, MBP, and EWAInit-BP under both parallel and serial scheduling, as illustrated in Fig. 7.
@@ IV.2.3 Planar surface code
Under parallel scheduling, the decoding logical error rates of LLRBP4 and MBP4 increase with increasing code length, while EWAInit-BP achieves a reduction in logical error rates at low physical error rates. Unfortunately, at higher error rates, decoding at higher code lengths becomes more challenging, preventing EWAInit-BP from achieving a reliable threshold on the planar surface code.
Under serial scheduling, EWAInit-BP performs better than MBP at $L=5$ and $L=7$; however, its performance at $L=9$ is not satisfying. In fact, the decoding failure rate of EWAInit-BP increases with higher code lengths, thus not fully demonstrating the error-correcting capabilities of the code. This is because, despite the dynamic updating of a priori probabilities, the method still partly relies on the initial input priors. We incorporate the OTS method as a ”post-processing” step for EWAInit, resetting the a priori probabilities of some nodes every few iterations, with the hyperparameters T=5 and C=20. This EWAInit-OTS method, while maintaining a time complexity of $O(n)$, shows improvements in convergence rate and logical error rates on the surface code compared to EWAInit-BP and other methods.
The patterns of average iterations of different algorithms for the planar surface code and the XZZX surface code are similar to those observed in the toric code and exhibit the same trends as the logical error rates. Therefore, we have omitted this data.
IV.2.4 XZZX Surface code
@@ V Conclusion
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we adopted optimization strategies from gradient descent in machine learning, resulting in Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update process and reduce the amplitude of oscillations. Simulations on trapping sets intuitively demonstrated the effectiveness of these algorithms.
@@
From the theoretical derivation of the message update rules of BP algorithm, we recognized the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm exhibits aggressive exploration and outperforms other algorithms without OSD post-processing on the toric code, planar surface code, and XZZX surface code. Specifically, EWAInit-BP shows high decoding accuracy even under parallel scheduling, with logical error rates decreasing as code length increases. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallelism while maintaining decoding accuracy.
However, the aggressive exploration and convergence capabilities of EWAInit-BP increase the probability of generating logical operators after error recovery. Therefore, directly applying this method to other error-correcting codes presents certain challenges. A potential research direction is to identify situations where logical operators may occur. Besides, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to more challenging noise models, potentially requiring specialized optimizations.
Acknowledgements.
This work was supported by the Colleges and Universities Stable Support Project of Shenzhen, China (No.GXWD20220817164856008) and Harbin Institute of Technology, Shenzhen - SpinQ quantum information Joint Research Center Project (No.HITSZ20230111). We would like to express our sincere gratitude to Zhengzhong Yi for his in-depth discussions on the principles of BP algorithms and meticulous feedback on the manuscript. We are also thankful to Zhipeng Liang for his assistance with the experimental code framework. Finally, we extend our appreciation to Prof. Xuan Wang for his insightful discussions and guidance on the gradient approach.

### v5

@@ IV-B2 Toric code
For the serial schedule, all three decoding algorithms demonstrated error correction capabilities. Among those, BP-OTS performs better than MBP at low physical error rates, but its performance is surpassed by MBP when the physical error rate exceeds 7%. The performance of EWAInit-BP, on the other hand, consistently outperforms these two methods across all tested conditions, with the logical error rate at $L=6$ being comparable to that of the other two algorithms at $L=8$, and the intersection of the logical error rate curves is approximately 15% [^2] We cannot yet consider the intersection of curves for short codes as the obtained threshold, as the convergence capability of EWAInit-BP degrades to some extent on larger codes. This will be discussed in Appendix VI-C.. In terms of iteration count, both EWAInit-BP and MBP converge in remarkably few iterations (often fewer than 5).
IV-B3 Planar surface code
The boundaries of the planar surface code are no longer contiguous, resulting in the emergence of some stabilizers with weight three and qubits that only connect two stabilizers at the boundaries. This configuration leads to a slight reduction in the error correction capabilities of the planar surface code compared to the Toric code. We compare the decoding logical error rates of LLR-BP4, MBP, and EWAInit-BP under both parallel and serial scheduling, as illustrated in Fig. 6.
@@ IV-B3 Planar surface code
Under parallel scheduling, the decoding logical error rates of LLRBP4 and MBP4 increase with increasing code length, whereas EWAInit-BP achieves a reduction in logical error rates at low physical error rates. Unfortunately, at higher error rates, decoding at higher code lengths becomes more challenging, preventing EWAInit-BP from achieving a reliable threshold on the planar surface code.
Under serial scheduling, EWAInit-BP performs better than MBP at $L=5$ and $L=7$; however, its performance at $L=9$ is not satisfying. Despite the dynamic updating of a priori probabilities, the method still partly relies on the symmetric initial input priors. One solution is to incorporate the OTS method as a ”post-processing” step for EWAInit, resetting the a priori probabilities of some nodes every few iterations, with the hyperparameters $T=5$ and $C=20$. This EWAInit-OTS method, while maintaining a time complexity of $O(n)$, shows improvements in convergence rate and logical error rates on the surface code compared to EWAInit-BP and other methods.
The patterns of average iterations of different algorithms for the planar surface code and the XZZX surface code are similar to those observed in the toric code and exhibit the same trends as the logical error rates. Therefore, we do not repeat the iteration figures here.
IV-B4 XZZX Surface code
@@
IV-C Generality on other quantum LDPC codes
Although the trapping set analysis presented in Section IV-A is difficult to directly extend to arbitrary quantum LDPC codes due to the structural differences among codes, the performance of our EWAInit-BP algorithm is not limited to 2D topological constructions. In this subsection, we provide empirical evidence that EWAInit-BP generalizes well to other classes of quantum LDPC codes. Two representative examples are considered: the Bivariate Bicycle codes proposed by IBM [50], and 3D Chamon codes which is an instance of 3D XYZ product of three repetition codes [51].
Fig. 9 presents a comparison of logical error rates between BP2-OSD-0 and EWAInit-BP under parallel scheduling on these two code families under depolarizing noise and code capacity noise model. For the Bivariate Bicycle Codes, EWAInit-BP almost consistently outperforms BP2-OSD-0 across all tested code distances, with its advantage more pronounced at lower physical error rates.
Similarly, for the 3D Chamon codes, EWAInit-BP demonstrates improved decoding performance for all tested distances. While both algorithms show typical threshold-like behavior, EWAInit-BP achieves lower logical error rates under comparable physical error rates. These results provide preliminary evidence that the exponential averaging mechanism is broadly beneficial for decoding quantum LDPC codes.
(a) Bivariate Bicycle Codes
(b) 3D Chamon Codes
Figure 9: Comparison of BP2-OSD-0 and EWAInit-BP on two non-surface-code quantum LDPC families: (a) Bivariate Bicycle Codes and (b) 3D Chamon Codes. EWAInit-BP almost outperforms BP-OSD-0 across all distances and physical error rates.
@@ V Conclusion
Since the a posteriori probability update of BP can be interpreted as the gradient optimization of an energy function, we adopted optimization strategies from gradient descent in machine learning, resulting in Momentum-BP and AdaGrad-BP. These mechanisms aim to smooth the update process and reduce the amplitude of oscillations. Simulations on trapping sets intuitively demonstrated the effectiveness of these algorithms, supporting the feasibility of leveraging gradient optimization principles for BP decoding.
@@
From the theoretical derivation of the message update rules of BP algorithm, we recognized the importance of a priori probabilities and transformed the message update smoothing into initial values adapting, proposing EWAInit-BP. This algorithm realizes an exponentially weighted average of past horizontal message updates and outperforms other algorithms without OSD post-processing on various surface codes. Specifically, EWAInit-BP shows high decoding accuracy even under parallel scheduling, with logical error rates decreasing as code length increases on short blocks. Notably, all our proposed methods do not introduce any additional factors to the time complexity of BP. Thus, EWAInit-BP can be implemented with $O(1)$ complexity in fully parallelism while maintaining decoding accuracy.
Furthermore, we also evaluated the performance of EWAInit-BP on two representative quantum LDPC codes—Bivariate Bicycle Codes and the 3D Chamon codes. Our algorithm demonstrates competitive decoding performance on both families under parallel scheduling, showing potential adaptability to high-dimensional or irregular topologies. The results suggest that the principle behind message adaptation strategies may extend beyond the specific context of surface codes, indicating broader theoretical significance. Future work will explore how to adapt our approach to other quantum LDPC codes with higher degeneracy or more complex structures, especially focusing on maintaining decoding performance at larger code distances.
However, applying exponential weighting to messages is still insufficient to ensure BP convergence for all code lengths. It may be necessary to draw inspiration from methods like neural BP to enhance the representational power of the algorithm. Besides, our simulations only considered Pauli errors on qubits and did not account for measurement errors, error propagation in circuits, etc. Future work needs to adapt BP to the circuit-level noise model, potentially requiring specialized optimizations.
VI Appendix
VI-A BP as gradient optimization
$J(\boldsymbol{Q})=$ $-\sum_{i=1}^{m}2\tanh^{-1}$
$\left((-1)^{s_{i}}\prod_{v_{j}\in\mathcal{N}(c_{i})}\tanh\left(\frac{\lambda_{H_{ij}}(Q_{v_{j}})}{2}\right)\right),$ (27)
where $\boldsymbol{Q}$ is the a posteriori probabilities of all the variable nodes in this ieration, $m$ is the number of check nodes. The term inside the parentheses measures the discrepancy between the result of the current iteration and the actual error syndrome of stabilizer $S_{i}$. The smaller the energy function, the closer the BP iteration result is to the correct solution.
The variables for gradient updates are the a posteriori estimates of each type of Pauli error of each variable node. Specifically, for the $W\in\{X,Y,Z\}$ type Pauli error on the $n$-th variable node, the a posteriori probability update is
$Q_{v_{j},W}^{(t)}=Q_{v_{j},W}^{(t-1)}-\alpha\frac{\partial J}{\partial Q_{v_{j},W}^{(t-1)}},$ (28)
where the partial derivative is
$\frac{\partial J}{\partial Q_{v_{j},W}}=$ $-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}\widetilde{m}_{c_{i}\rightarrow v_{j}}$
$*\frac{g_{ij}(Q)\ e^{-Q_{v_{j},W}}}{e^{-Q_{v_{j},X}}+e^{-Q_{v_{j},Y}}+e^{-Q_{v_{j},Z}}-e^{-Q_{v_{j},H_{ij}}}},$ (29)
where $g_{ij}$ is a term that is always greater than 0 (see [21]), and $\widetilde{m}_{c_{i}\rightarrow v_{j}}$ is the a posteriori probability summation term, approximated using the $\tanh$ function
$\widetilde{m}_{c_{i}\rightarrow v_{j}}=(-1)^{s_{i}}\prod_{v_{j^{\prime}}\in\mathcal{N}(c_{i})\backslash v_{j}}\tanh\left(\frac{\lambda_{H_{ij^{\prime}}}(Q_{v_{j^{\prime}}})}{2}\right).$ (30)
If we view the last term on the RHS of (29) as a variable coefficient, then the posterior update in BP can be seen as a gradient descent step that omits this coefficient term.
VI-B Iterative expansion of EWA-based BP algorithms
Here we present the expansion of Momentum-BP and EWAInit-BP proposed in this article to clarify their properties. To streamline calculations and maintain consistency, we set $\gamma$ in Momentum-BP to 0.
VI-B1 Momentum-BP
For Momentum-BP, the a posterori probabilities are updated in the first iteration as
$Q_{v_{j},W}^{(1)}$ $=\alpha(\Pi_{v_{j},W}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(1)})+(1-\alpha)\Pi_{v_{j},W}^{(0)}$ (31)
$=\Pi_{v_{j},W}^{(0)}+\alpha\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(1)},$ (32)
where the first term on the RHS represents the calculation in the current iteration. The second term represents the calculation from the pervious iteration, which in the first iteration is simply the a priori probability, say $Q_{v_{j},W}^{(0)}=\Pi_{v_{j},W}^{(0)}$.
Now we can expand the a posterori update iteratively,
$Q_{v_{j},W}^{(t)}$ $=\alpha(\Pi_{v_{j},W}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)})+(1-\alpha)Q_{v_{j},W}^{(t-1)}$ (33)
$=\alpha(\Pi_{v_{j},W}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)})+(1-\alpha)*$
$\quad(\alpha(\Pi_{v_{j},W}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-1)})+(1-\alpha)Q_{v_{j},W}^{(t-2)})$ (34)
$=\alpha\sum_{k=0}^{t-1}(1-\alpha)^{k}\left(\Pi_{v_{j},W}^{(0)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)}\right)$
$\quad+(1-\alpha)^{t}Q_{v_{j},W}^{(0)}.$ (35)
Using the initial condition $Q_{v_{j},W}^{(0)}=\Pi_{v_{j},W}^{(0)}$, we can substitute $Q_{v_{j},W}^{(0)}$ in (35) to obtain
$Q_{v_{j},W}^{(t)}$ $=\Pi_{v_{j},W}^{(0)}\left(\alpha\sum_{k=0}^{t-1}(1-\alpha)^{k}+(1-\alpha)^{t}\right)$
$\quad+\alpha\sum_{k=0}^{t-1}(1-\alpha)^{k}\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)}.$ (36)
Now, let’s examine the coefficient of $\Pi_{v_{j},W}^{(0)}$:
$\alpha\sum_{k=0}^{t-1}(1-\alpha)^{k}+(1-\alpha)^{t}.$ (37)
This is a finite geometric series, and we can simplify it using the formula
$\sum_{k=0}^{n}r^{k}=\frac{1-r^{n+1}}{1-r}.$ (38)
Substituting back to (37), we get
$\alpha\sum_{k=0}^{t-1}(1-\alpha)^{k}$ $+(1-\alpha)^{t}=$ (39)
$\alpha\cdot\frac{1-(1-\alpha)^{t}}{\alpha}+(1-\alpha)^{t}=1.$ (40)
Thus, we have shown that the coefficient of $\Pi_{v_{j},W}^{(0)}$ is indeed 1. The complete general term formula is therefore
$Q_{v_{j},W}^{(t)}=\Pi_{v_{j},W}^{(0)}+\sum_{k=0}^{t}\alpha(1-\alpha)^{k}\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)}.$ (41)
VI-B2 EWAInit-BP
Similarly, for EWAInit-BP, we have the initial condition
$\Pi_{v_{j},W}^{(1)}$ $=\Pi_{v_{j},W}^{(0)},$ (42)
$Q_{v_{j},W}^{(1)}$ $=\Pi_{v_{j},W}^{(1)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(1)},$ (43)
and the general term form
$\Pi_{v_{j},W}^{(t)}$ $=\alpha\Pi_{v_{j},W}^{(0)}+(1-\alpha)Q_{v_{j},W}^{(t-1)}$ (44)
$=\Pi_{v_{j},W}^{(0)}+\sum_{k=1}^{t}(1-\alpha)^{k}\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)},$ (45)
$Q_{v_{j},W}^{(t)}$ $=\Pi_{v_{j},W}^{(t)}+\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)}$ (46)
$=\Pi_{v_{j},W}^{(0)}+\sum_{k=0}^{t}(1-\alpha)^{k}\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)},$ (47)
where the coefficient of $\Pi_{v_{j},W}^{(0)}$ is also 1, derived by the same approach.
Although the approaches for improvement are different, both Momentum-BP and EWAInit-BP attempt to enhance the update of the a posteriori probability $Q_{v_{j},W}^{(t)}$ by applying a weighted combination of past messages. The distinction lies in the fact that the recurrence relation derived from performing EWA on the prior probability is more concise and intuitive, with only a $(1-\alpha)$ factor applied to past messages. Interestingly, EWAInit-BP, with its simpler recurrence relation, also demonstrates better performance in simulations.
VI-C Compared to Adaptive MBP
Adaptive MBP (AMBP)[21] defines a parameter range (for surface codes, $\alpha\in[1,0.5]$) and iterates through the possible values from large to small until MBP converges with a specific parameter $\alpha^{*}$, at which point it returns the result. Sorting the parameters from large to small prevents overly aggressive message updates that might cause logical errors, while ensuring convergence in most cases. In this appendix, we uniformly sampled 11 values within the range $[1,0.5]$, making the worst-case time overhead for AMBP approximately 10 times that of MBP. We can utilize a similar adaptive parameter selection strategy for EWAInit-BP using a parameter range of $\alpha\in[1,0]$, where the parameter represents the strength of the EWA. We refer to the resulting algorithm as AEWA-BP (Adaptive EWAInit-BP).
We compare EWAInit-BP, AEWA-BP with parallel scheduling, and AMBP with both parallel and serial scheduling on planar surface codes of sizes $[[13,1,3]]$, $[[85,1,7]]$, and $[[221,1,11]]$, as shown in Fig. 10. EWAInit-BP and parallel AMBP no longer exhibit the property of decreasing logical error rates with increasing code length, indicating that large surface codes remain a challenge for parallel BP. However, AEWA-BP with parallel scheduling demonstrates accuracy comparable to AMBP with serial scheduling. A major issue with AEWA-BP is that its parameter range cannot cover all possible magnitudes of message updates, meaning it theoretically cannot guarantee convergence for all errors.
Figure 10: Performance of EWAInit-BP, AEWA-BP, parallel AMBP, and serial AMBP on the planar surface code with L = 3, 7, 11. EWAInit-BP and parallel AMBP achieve similar accuracy, but both exhibit higher error rates at $L=11$ compared to $L=7$. AEWA-BP achieves accuracy comparable to serial AMBP.
Next, we compared the runtime of the aforementioned algorithms to verify the high overhead of adaptive methods discussed in the main text, as shown in Fig. 11. The runtime of parallel AMBP is very similar to that of AEWA-BP and is therefore omitted from the figure. As shown, EWAInit-BP exhibits the same scaling properties as traditional BP: its runtime grows linearly with code length and increases with physical error rates due to the rising number of iterations (up to a maximum of 100).
Figure 11: Average running time of EWAInit-BP, AEWA-BP, and serial AMBP on the planar surface code with L = 3, 7, 11. For the latter two algorithms that utilize adaptive methods, the running time increases significantly as the physical error rate grows.
However, algorithms utilizing adaptive methods show higher scaling. As the physical error rate increases, these algorithms tend to fully traverse the parameter range, resulting in runtime scaling not only with the number of iterations but also with the parameter range length in high-error-rate scenarios. In this experiment, the parameter range is set to 11 values. It can be observed that the runtime of adaptive methods is an order of magnitude higher than that of EWAInit-BP under high error rates. Table. II presents the runtime (ms) of the three algorithms under several error rates, offering a more intuitive perspective.
TABLE II: Comparison of the running time of EWAInit-BP, AEWA-BP, and AMBP under certain physical error rates
PER EWAInit-BP (ms) AEWA-BP (ms) AMBP (ms)
$10^{-2}$ 1.1 1.6 1.6
$6\times 10^{-2}$ 3.7 27.9 41.3
$10^{-1}$ 7.7 81.9 137.0
$2\times 10^{-1}$ 20.1 240.6 405.6